A worker received a $10,000 bonus and decided to split it among three different accounts. He placed part in a savings account paying 4.5% per year, twice as much in government bonds paying 5%, and the rest in a mutual fund that returned 4%. His income from these investments after one year was $455. How much did the worker place in each account?
Answer by mananth(16946)

$ 10,1000 - Total
Savings = x------- 4.5%
Bonds=2x-------5%
(10000-3x)-----4%
Interest = 455
4.5%x+5%*(2x)+(10000-3x)*4%= 455
multiply equation by 100
4.5x+10x+40000-12x=45500
2.5x=45500-40000
2.5x=5500
x= 5500/2.5
x= 2200
so you know how much he invested in other accounts